INAF queue observing, 2017 Nov 26-27

Observer: F. Cusano, A. Gargiulo
Telescope Operator: S. Allanson
Support Astronomer: D. Thompson
Instrument: LUCI

Summary

At the sunset the sky is clear.
LUCI1 has still problems so we can not use it, the ad-sec dx is ok and we can use also for AO.

Night duration (12\xB0 el twilight) 1:40am- 7:06am UT = 11 h

Observing time: * *10h 10*(%)*

Weather time loss: * *h (%)

technical time loss * *1h (%)

Data Summary

Project_RUN Objects Instrument Filter ExpTime Notes
Antoniucci V1057Cyg LUCI zJ, HK, K (48, 10, 10)s no imaging
Antoniucci V2493Cyg LUCI zJ, HK (960, 960)s  
Antoniucci V733cep LUCI zJ, HK (32, 30)s no imaging
Antoniucci RwAur LUCI zJ, HK (56, 28)s no imaging
Antoniucci DGTau LUCI AO   1h  
Vanzella jDz11 LUCI AO Ks 40 min  
Grazian COSMOS308 MODS dual grating 1.75h x2(bino) completed
Standard Ffeige34 MODS dual grating    
Calibration Arc, lamp, slit 0.75" LUCI      

Details

00:50 starting with LUCI2 only

00:55 slewing to Antonniucci Telluric of V1057Gyg
  • acq luci2.20171127.0032-34
  • science zJ filter luci2.20171127.0035-38
  • science HK filter luci2.20171127.0039-42
  • science K filter luci2.20171127.0043-46
  • seeing 0.7" on guider at 1:06

01:10 slewing to V1057Cyg
  • acq luci2.20171127.0047-53
  • science zJ filter luci2.20171127.0054-57
  • science HK filter luci2.20171127.0058-61
  • science K filter luci2.20171127.0062-65
  • seeing 0.5" on guider at 1:22

01:38 slewing to Telluric of V2493Cyg
  • acq luci2.20171127.0066-68
  • science zJ filter luci2.20171127.0069-72=
  • science HK filter luci2.20171127.0073-76
  • seeing 0.6" on guider at 1:45

01:49 slewing to V2493Cyg
  • acq luci2.20171127.0077-80
  • science zJ filter luci2.20171127.0081-84
  • science HK filter luci2.20171127.0085-88
  • seeing 0.5" on guider at 2:22
  • imaging luci2.20171127.0089-93

02:50 telluric for Nisini Field2
  • acq luci2.20171127.0094-96
  • science J filter luci2.20171127.0097-100
  • science H filter luci2.20171127.0101-104
  • science K filter luci2.20171127.0105-108
  • seeing 0.7 on guider at 2:52

03:21 slewing to Field 2 of Nisini
  • problem in aligning the mask, the lms file does not match luci2.20171127.0109-111
  • we try to resolve the problem, but we did not succeded, D. Thompson suggest to change object
  • error in the Nisini program is not easily solvable, we move to Antoniucci targets
  • we will update the log soon

03:45 slewing to 733Cep of Antoniucci
  • seeing 0.5 " on guider at 4:00
  • acq luci2.20171127.0112-116
  • science zJ filter luci2.20171127.0117-120
  • science HK filter luci2.20171127.0121

*04:07 telluric for 733Cep
  • seeing 0.6" on the guider at 4:20
  • acq ==luci2.20171127.0122-124 =
  • science zJ filter luci2.20171127.0125-128
  • science HK filter luci2.201711.0129-0132

04:21 slewing to RWAur of Antoniucci
  • acq luci2.20171127.0133-136
  • sci zJ filter luci2.20171127.0137-140
  • science HK filter luci2.201711.0141-0144

04:38 telluric for RWAur of Antoniucci
  • seeing 0.6" on guider at 4:48
  • acq luci2.20171127.0145-147
  • sci zJ filter luci2.20171127.0148-151
  • science HK filter luci2.201711.0152-0155

05:00 slewing to LUCI2 AO for DGTau of Antoniucci
  • seeing 0.6" on guider at 5:05
  • science luci2.20171127.0156-213
  • dimm 0.9" at 7:05

07:10 the AO corrections do not appear as good as expected. We stop the run. Doug, Enrico and Guido check the possible reasons.

07:19 we continue with the observations
  • dimm 0.9" at 7:23

07:45 AO loop stops. The AO-SAO works to fix the problem.

07:58 slewing to the PSF star.

08:03 slewing to another PSF star: the previous was too faint.

08:17 acquisition start. To close the loop takes long time.

08:26 slewing to target jDz11 of Vanzella
  • binning 3x3, NCPA OFF
  • dimm 1.0" at 8:49
  • observations in pause: opened loop at 8:51
  • 8:53 start with the observations

09:19 no flexure corrections due to a problem with a software. We stop the observations. The SAO and the TO check the impact on the images and it results to not be a big issue.

09:32 start the observation again.
  • dimm 1.0" at 9:33

09:50 seeing limited observations of Vanzella

10:02 slewing to MODS

10:24 slewing to Field_I of Grazian
  • some high cirrus coming in at 10:17
  • seeing on guider 0.7", dimm seeing 0.6", 10:28
  • acqBinoMODs mods1r.20171127.0002-004 ; mods2r.20171127.0002-004

10:36 execBinoMODS pointI_mos_4_meta.obs
  • mods1r.20171127.0005, mods1b.20171127.0002 ; mods2r.20171127.0005, mods2b.20171127.0002
  • seeing 0.6" on guider, dimm 0.6"
  • we see some high cirrus on the all sky cam

10:56 execBinoMODS pointI_mos_5.obs
  • mods1r.20171127.0006-7, mods1b.20171127.0003-4 ; mods2r.20171127.0006-7, mods2b.20171127.0003-4

11:32 execBinoMODS pointI_mos_6.obs
  • mods1r.20171127.0008-9, mods1b.20171127.0005-6 ; mods2r.20171127.0008-9, mods2b.20171127.0005-6
  • seeing 0.6" on guider, dimm 0.56" at 11:33

12:09 execBinoMODS pointI_mos_7.obs
  • mods1r.20171127.00010-11, mods1b.20171127.0007-8 ; mods2r.20171127.00010-11, mods2b.20171127.0007-8
  • seeing 0.5" on guider, dimm 0.6" at 12:09

12:45 slewing to feige34
  • seeing 0.5 on guider, dimm at 12:45
  • mods1r.20171127.00012-13 ; mods2r.20171127.00012-13

12:59 execBinoMODS feige 34
  • mods1r.20171127.00014-16, mods1b.20171127.0009-11 ; mods2r.20171127.00014-16, mods2b.20171127.0009-11

*13:25 * slewing to LUCI calibration
  • flat(bias) - zJ G200 =luci2.20171127.0303-312
  • arc =luci2.20171127.0313-316
  • flat(bias) - HK G200 luci2.20171127.0317-326
  • arc =luci2.20171127.0327-332
  • flat(bias) - K G210 luci2.20171127.0333-342
  • arc luci2.20171127.0343-348
Topic revision: r9 - 06 Jul 2018, FeliceCusano
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback